# A motivic Fundamental Lemma

### v3

@@ 4.8. Geometric Stabilization
We fix a field $k$ of characteristic zero, a smooth projective curve $X$ of genus $g$ over $k$ with a rational point $\infty\in X(k)$, an even integer $d$, larger or equal to $22g-$ 2 and an integer $r$. We view $X$ embedded in projective space as in Section 4.7. We consider an ample line bundle $D$ on $X$ of degree $d$, given by trivializations on open charts. As in Section 4.6, we consider a definable quasi-split reductive group ${\mathbf{G}}$ over $X$, with parameters $b,\tau,f$.
@@
The homomorphism $\kappa:\hat{\mu}\to\widehat{\mathbb{T}}$ corresponds to an element in ${\mathbb{X}}_{*}(\widehat{\mathbb{T}})\otimes{\mathbb{Q}}/{\mathbb{Z}}=\operatorname{\mathsf{Hom}}({\mathbb{X}}_{*}({\mathbb{T}}),{\mathbb{Q}}/{\mathbb{Z}})$. We consider $\kappa$ as a character of ${\mathbb{X}}_{*}({\mathbb{T}})$ by identifying ${\mathbb{Q}}/{\mathbb{Z}}$ with the roots of unity in $\Lambda$ via ${\mathbb{Q}}/{\mathbb{Z}}\xrightarrow{e^{2\pi i(\cdot)}}\Lambda$. If follows from [37, Sections 6.2.2 & 6.3] that $\kappa$ factors through $\pi_{0}(\widetilde{\mathcal{P}}^{{\mathrm{ani}}}_{G})$ when restricted to ${\widetilde{{\mathbf{A}}}}_{{\mathcal{E}}}^{\mathrm{ani}}$. Thus we obtain a well-defined $\kappa$-isotypical component $\chi_{\mathrm{Psf},\widetilde{{\mathbf{A}}}^{{\mathrm{ani}}}_{H_{\mathcal{E}}},\mathrm{rel}}(\widetilde{{\mathbf{M}}}_{G{|\widetilde{{\mathbf{A}}}^{{\mathrm{ani}}}_{H_{\mathcal{E}}}}},{\kappa})$ in the sense of Section 4.4.
@@
For every quasi-split form of ${\mathbb{G}}$ over $X_{\mathbb{F}}$, $G\to X_{\mathbb{F}}$, that splits over a Galois cover of degree $r$, the specialization of $\widetilde{{\mathbf{M}}}_{G}$ to ${\mathbb{F}}$ for some choice of parameters is the set of ${\mathbb{F}}$-points of $\widetilde{M}_{G}$, the coarse moduli space of ~$\widetilde{{\mathcal{M}}}^{\mathrm{ani}}_{G}$. The same holds for the various other geometric object constructed above, such as ${\mathbf{P}}_{G}$ and ${\mathbf{A}}$.
@@
since this is the way we defined the $\kappa$-isotypical components for motives. By [22, Lemma 6.6], this sum is equal to #$\#^{\kappa}\widetilde{{\mathcal{M}}}_{G,a}({\mathbb{F}})$. ∎
Recall the Geometric Stabilization theorem of Ngô [37, Theorem 6.4.2], as well as the geometric version of Waldspurger non-standard Fundamental Lemma [37, Theorem 8.8.2]. See also [22, Theorems 1.1 and 1.3].
@@
It suffices to combine Proposition 4.8.4 with the motivic statements, Theorems 4.8.1 and 4.8.2. ∎
@@ 5. Proof of the main theorems
In this section we prove Theorems 4.8.1 and 4.8.2 by reducing them gradually to a duality statement about motivic integrals on generic Hitchin fibers, Theorem 5.4.1.
@@ 5.1. Reduction to a point
By Proposition 2.6.2, Theorems 4.8.1 and 4.8.2 follow from their following fiberwise versions.
@@ 5.2. Reduction to inertia stacks
For now on, we fix a pseudo-finite field $K$ and $a\in\widetilde{A}^{{\mathrm{ani}}}(K)$. Then $\pi_{0}(\widetilde{{\mathcal{P}}}_{G,a})$ is a finite constant group scheme by [22, Prop. 4.38].
For $b$ generic, $\widetilde{{\mathcal{P}}}_{G,b}$ is a proper commutative group scheme, with group of connected components isomorphic to the center $Z(X,\widehat{G})$ of $\widehat{G}$. Recall from Section 4.6 that we consider the group of characters ${\mathbb{X}}^{*}(\widehat{\mathbb{T}})$ via a basis of this lattice, which is a definable set. Since $\widehat{\mathbb{T}}/Z(X,\widehat{G})$ is also a torus, we also have a finite definable set consisting of a basis of ${\mathbb{X}}^{*}(\widehat{\mathbb{T}}/Z(X,\widehat{G}))$. Let $\lambda$ be the composition of the inclusion ${\mathbb{X}}^{*}(\widehat{\mathbb{T}}/Z(X,\widehat{G}))\subseteq{\mathbb{X}}^{*}(\widehat{\mathbb{T}})$ with the natural surjection $\widetilde{A}^{{\mathrm{ani}}}\times{\mathbb{X}}^{*}(\widehat{\mathbb{T}})={\mathbb{X}}_{*}({\mathbb{T}})\to\pi_{0}(\widetilde{\mathcal{P}}_{G})$ from [22, Prop. 4.38] relative over $\widetilde{A}^{{\mathrm{ani}}}$. Let $\square$ be the image of $\lambda$ in $\pi_{0}(\widetilde{\mathcal{P}}_{G})$. For any $a\in\widetilde{A}^{{\mathrm{ani}}}(K)$ the fiber $\square_{a}$ is definable, since it is generated by the image of the basis of ${\mathbb{X}}^{*}(\widehat{\mathbb{T}}/Z(X,\widehat{G}))$ that is definable.
Hence for generic $b$, $\square_{b}$ is the trivial group, and for every $a\in\widetilde{A}^{{\mathrm{ani}}}(K)$, $\square_{a}$ is a finite constant group scheme. We then define by fiber product $\widetilde{\mathcal{P}}^{\square}_{G}=\square\times_{\pi_{0}(\widetilde{\mathcal{P}}_{G})}\widetilde{\mathcal{P}}_{G}$, and similarly ~$\widetilde{\mathcal{M}}^{\square}_{G}$. The coarse moduli spaces of ~$\widetilde{\mathcal{M}}^{\square}_{G}$ and $\widetilde{\mathcal{P}}^{\square}_{G}$ can be seen as definable sets denoted by $\widetilde{{\mathbf{M}}}^{\square}_{G}$ and $\widetilde{{\mathbf{P}}}^{\square}_{G}$ respectively.
@@
~$\widetilde{\mathcal{M}}^{\square,t}_{G,a}:=\widetilde{\mathcal{M}}^{\square}_{G,a}\times^{\widetilde{{\mathcal{P}}}^{\square}_{G,a}}T_{t}.$
@@
First define the substack ~$\widetilde{{\mathcal{M}}}^{\lozenge}_{H_{\mathcal{E}}}\subseteq\widetilde{{\mathcal{M}}}_{H_{\mathcal{E}}}^{G}$ as the maximal open substack such that the following diagram commutes
@@
As in Section 4.7, $\widetilde{{\mathcal{P}}}_{H_{\mathcal{E}}}$ embeds into ~$\widetilde{{\mathcal{M}}}_{H_{\mathcal{E}}}$ by means of the Kostant section and we define $\widetilde{{\mathcal{P}}}^{\lozenge}_{H_{\mathcal{E}}}$ as the intersection of $\widetilde{{\mathcal{P}}}_{H_{\mathcal{E}}}$ with ~$\widetilde{{\mathcal{M}}}^{\lozenge}_{H_{\mathcal{E}}}$. As before, we get associated definable sets $\widetilde{{\mathbf{M}}}^{\lozenge}_{H_{\mathcal{E}}}$ and $\widetilde{{\mathbf{P}}}^{\lozenge}_{H_{\mathcal{E}}}$.
There is a canonical morphism $\pi_{0}(\widetilde{{\mathcal{P}}}^{\square}_{G,a})\to\pi_{0}(\widetilde{{\mathcal{P}}}^{\lozenge}_{H_{\mathcal{E}},a})$, allowing to define for $t\in\pi_{0}(\widetilde{{\mathcal{P}}}^{\square}_{G,a})$ twists ~$\widetilde{{\mathcal{M}}}^{\lozenge,t}_{H_{\mathcal{E}},a}$, with associated definable $\widetilde{{\mathbf{M}}}^{\lozenge,t}_{H_{\mathcal{E}},a}$, which are compatible with twists of $I_{\hat{\mu}}\widetilde{{\mathcal{M}}}_{G}^{\square}$ in the following sense.
@@
indexed by orbits of definable coendoscopic data ${\mathcal{E}}=(\tau,\rho_{\tau},\rho_{\tau}\to\rho)$ under the natural $Z(X,G)$-action. Moreover, assume that there is a $k$-rational point $\infty_{\rho_{\tau}}$ of $\rho_{\tau}$ mapping to $\infty_{\rho}$ under $\rho_{\tau}\to\rho$. Then the morphism $\overline{\mu}_{{\mathcal{E}}}$ from (3.6.3) induces a definable bijection $\widetilde{{\mathbf{M}}}^{\lozenge,t}_{H_{\mathcal{E}},a}\cong I_{\hat{\mu}}\widetilde{{\mathbf{M}}}^{\square,t}_{G,a,{\mathcal{E}}}$.
@@
Moreover, if ${\mathcal{E}}$ is such that there is a $k$-rational point $\infty_{\rho_{\tau}}$ of $\rho_{\tau}$ mapping to $\infty_{\rho}$ under $\rho_{\tau}\to\rho$, then the construction (3.6.3) can be twisted by $t$, yielding a definable map
$\widetilde{{\mathbf{M}}}^{\lozenge,t}_{H_{\mathcal{E}},a}\longrightarrow I_{\hat{\mu}}\widetilde{{\mathbf{M}}}^{\square,t}_{G,a,{\mathcal{E}}}.$
This map is a bijection in every finite field by [22, Lemma 6.8], hence in every pseudo-finite field. ∎
@@
Given $t\in{\mathbb{X}}^{*}(\widehat{\mathbb{T}}/Z(X,\widehat{G}))$, we have canonical maps
@@
We also have the map $\lambda_{a}\colon{\mathbb{X}}_{*}({\mathbb{T}}/Z(X,G))\to\pi_{0}(\widetilde{{\mathcal{P}}}^{\square}_{G,a})$. We write abusively also $t$ for the image of $t$ in $\pi_{0}(\widetilde{{\mathcal{P}}}^{\square}_{G,a})$ through this composition of maps. In particular, given $t\in{\mathbb{X}}^{*}(\widehat{\mathbb{T}}/Z(X,\widehat{G}))$, we write ~$\widetilde{{\mathcal{M}}}^{\square,t}_{G,a}$ for the twist of ~$\widetilde{{\mathcal{M}}}^{\square}_{G,a}$ by the image of $t$ in $\pi_{0}(\widetilde{{\mathcal{P}}}^{\square}_{G,a})$.
@@
For every pseudo-finite field $K$, $a\in\widetilde{{\mathbf{A}}}^{{\mathrm{ani}}}(K)$, $t\in{\mathbb{X}}^{*}(\widehat{\mathbb{T}}/Z(X,\widehat{G}))$ and $s\in{\mathbb{X}}^{*}({\mathbb{T}}/Z(X,G))$, we have
@@
$\chi_{\mathrm{Psf},\mathrm{rel}}(I_{\hat{\mu}}\widetilde{{\mathbf{M}}}_{\widehat{G},a,{\mathcal{E}}_{H}}^{\square,s})=\chi_{\mathrm{Psf},\mathrm{rel}}(\widetilde{{\mathbf{M}}}_{H,a}^{\lozenge,s}).$
@@
$\sum_{s,t}\chi_{\mathrm{Psf},\mathrm{rel}}(\widetilde{{\mathbf{M}}}_{H,a}^{\square,s})=\left\lvert\pi_{0}(\widetilde{{\mathbf{P}}}^{\square}_{G,a})\right\rvert\left\lvert\pi_{0}(\widetilde{{\mathcal{P}}}^{\square}_{\widehat{G},a})\right\rvert\chi_{\mathrm{Psf},\mathrm{rel}}(\widetilde{{\mathbf{M}}}_{H,a}^{\lozenge},\mathrm{stab}).$
From Lemma 6.10 in [22], $\dim(\widetilde{\mathbf{M}}_{G})-w(\kappa_{a})=r_{G}^{H}(D)$, so we now conclude using Lemma 4.4.3. ∎
@@ 5.3. Reduction to motivic integrals
The restrictions of ${\mathcal{M}}_{G}$ and ${\mathcal{P}}_{G}$ to $U_{a}$ are denoted by ${\mathcal{M}}_{G,U_{a}}$ and ${\mathcal{P}}_{G,U_{a}}$. Essentially by definition of Henselization, see for example [36, Theorem I.4.2], we have
@@
For $t\in\pi_{0}(\widetilde{{\mathcal{P}}}^{\square}_{G,a})$ we then get a smooth DM-stack ~$\widetilde{{\mathcal{M}}}^{\square,t}_{G,U_{a}}$ and a corresponding definable $\widetilde{{\mathbf{M}}}^{\square,t}_{G,B_{a}}$.
@@
Let $K$ be a pseudo-finite field $K$. For every $a\in\widetilde{{\mathbf{A}}}^{{\mathrm{ani}}}(K)$, every $t\in{\mathbb{X}}^{*}(\widehat{\mathbb{T}}/Z(X,\widehat{G}))$ and every $s\in{\mathbb{X}}^{*}({\mathbb{T}}/Z(X,G))$, we have

### v4

@@ 4.9. Geometric Stabilization
We fix a field $k$ of characteristic zero, a smooth projective curve $X$ of genus $g$ over $k$ with a rational point $\infty\in X(k)$, an even integer $d$, larger or equal to 2g22g-2 and an integer $r$. We view $X$ embedded in projective space as in Section 4.7. We consider an ample line bundle $D$ on $X$ of degree $d$, given by trivializations on open charts. As in Section 4.6, we consider a definable quasi-split reductive group $\mathbf{G}$} over $X$, with parameters $b,\tau,f$.
@@
The homomorphism $\kappa:\hat{\mu}\to\widehat{\mathbb{T}}$ corresponds to an element in ${\mathbb{X}}_{*}(\widehat{\mathbb{T}})\otimes{\mathbb{Q}}/{\mathbb{Z}}=\operatorname{\mathsf{Hom}}({\mathbb{X}}_{*}({\mathbb{T}}),{\mathbb{Q}}/{\mathbb{Z}})$. We consider $\kappa$ as a character of ${\mathbb{X}}_{*}({\mathbb{T}})$ by identifying $\mathbb{Q}$$\mathbb{Z}$} with the roots of unity in $\Lambda$ via ${\mathbb{Q}}/{\mathbb{Z}}\xrightarrow{e^{2\pi i(\cdot)}}\Lambda$. If follows from [38, Sections 6.2.2 & 6.3] that $\kappa$ factors through $\pi_{0}(\widetilde{\mathcal{P}}^{{\mathrm{ani}}}_{G})$ when restricted to ${\widetilde{{\mathbf{A}}}}_{{\mathcal{E}}}^{\mathrm{ani}}$. Thus we obtain a well-defined $\kappa$-isotypical component $\chi_{\mathrm{Psf},\widetilde{{\mathbf{A}}}^{{\mathrm{ani}}}_{H_{\mathcal{E}}},\mathrm{rel}}(\widetilde{{\mathbf{M}}}_{G{|\widetilde{{\mathbf{A}}}^{{\mathrm{ani}}}_{H_{\mathcal{E}}}}},{\kappa})$ in the sense of Section 4.4.
@@
For every quasi-split form of $\mathbb{G}$} over $X_{\mathbb{F}}$, $G\to X_{\mathbb{F}}$, that splits over a Galois cover of degree $r$, the specialization of $\widetilde{{\mathbf{M}}}_{G}$ to $\mathbb{F}$} for some choice of parameters is the set of $\mathbb{F}$}-points of $\widetilde{M}_{G}$, the coarse moduli space of $\widetilde{{\mathcal{M}}}^{\mathrm{ani}}_{G}$. The same holds for the various other geometric object constructed above, such as ${\mathbf{P}}_{G}$ and $\mathbf{A}$}.
@@
since this is the way we defined the $\kappa$-isotypical components for motives. By [23, Lemma 6.6], this sum is equal to $^{\kappa}\widetilde{{\mathcal{M}}}_{G,a}({\mathbb{F}})$. ∎
Recall the Geometric Stabilization theorem of Ngô [38, Theorem 6.4.2], as well as the geometric version of Waldspurger non-standard Fundamental Lemma [38, Theorem 8.8.2]. See also [23, Theorems 1.1 and 1.3].
@@
It suffices to combine Proposition 4.9.4 with the motivic statements, Theorems 4.9.1 and 4.9.2. ∎
@@ 5. Proof of the main theorems
In this section we prove Theorems 4.9.1 and 4.9.2 by reducing them gradually to a duality statement about motivic integrals on generic Hitchin fibers, Theorem 5.4.1.
@@ 5.1. Reduction to a point
By Proposition 2.6.2, Theorems 4.9.1 and 4.9.2 follow from their following fiberwise versions.
@@ 5.2. Reduction to inertia stacks
For now on, we fix a pseudo-finite field $K$ and $a\in\widetilde{A}^{{\mathrm{ani}}}(K)$. Then $\pi_{0}(\widetilde{{\mathcal{P}}}_{G,a})$ is a finite constant group scheme by [23, Prop. 4.38].
[…]
Hence for generic $b$, $\square_{b}$ is the trivial group, and for every $\in\widetilde{A}^{{\mathrm{ani}}}(K), a\square_{a}$ is a finite constant group scheme. We then define by fiber product $\widetilde{\mathcal{P}}^{\square}_{G}=\square\times_{\pi_{0}(\widetilde{\mathcal{P}}_{G})}\widetilde{\mathcal{P}}_{G}$, and similarly $\widetilde{\mathcal{M}}^{\square}_{G}$. The coarse moduli spaces of $\widetilde{\mathcal{M}}^{\square}_{G}$ and $\widetilde{\mathcal{P}}^{\square}_{G}$ can be seen as definable sets denoted by $\widetilde{{\mathbf{M}}}^{\square}_{G}$ and $\widetilde{{\mathbf{P}}}^{\square}_{G}$ respectively.
@@
$\widetilde{\mathcal{M}}^{\square,t}_{G,a}:=\widetilde{\mathcal{M}}^{\square}_{G,a}\times^{\widetilde{{\mathcal{P}}}^{\square}_{G,a}}T_{t}.$
@@
[…]
As in Section 4.7, $\widetilde{{\mathcal{P}}}_{H_{\mathcal{E}}}$ embeds into $\widetilde{{\mathcal{M}}}_{H_{\mathcal{E}}}$ by means of the Kostant section and we define $\widetilde{{\mathcal{P}}}^{\lozenge}_{H_{\mathcal{E}}}$ as the intersection of $\widetilde{{\mathcal{P}}}_{H_{\mathcal{E}}}$ with $\widetilde{{\mathcal{M}}}^{\lozenge}_{H_{\mathcal{E}}}$. As before, we get associated definable sets $\widetilde{{\mathbf{M}}}^{\lozenge}_{{\mathcal{E}}}$ and $\widetilde{{\mathbf{P}}}^{\lozenge}_{H_{\mathcal{E}}}$.
There is a canonical morphism $\pi_{0}(\widetilde{{\mathcal{P}}}^{\square}_{G,a})\to\pi_{0}(\widetilde{{\mathcal{P}}}^{\lozenge}_{H_{\mathcal{E}},a})$, allowing to define for $t\in\pi_{0}(\widetilde{{\mathcal{P}}}^{\square}_{G,a})$ twists $\widetilde{{\mathcal{M}}}^{\lozenge,t}_{{\mathcal{E}},a}$, with associated definable $\widetilde{{\mathbf{M}}}^{\lozenge,t}_{{\mathcal{E}},a}$, which are compatible with twists of $I_{\hat{\mu}}\widetilde{{\mathcal{M}}}_{G}^{\square}$ in the following sense.
@@
indexed by orbits of definable coendoscopic data ${\mathcal{E}}=(\tau,\rho_{\tau},\rho_{\tau}\to\rho)$ under the natural $Z(X,G)$-action. Moreover, the morphism $\overline{\mu}_{{\mathcal{E}}}$ from (3.6.3) induces a definable bijection $\widetilde{{\mathbf{M}}}^{\lozenge,t}_{{\mathcal{E}},a}\cong I_{\hat{\mu}}\widetilde{{\mathbf{M}}}^{\square,t}_{G,a,{\mathcal{E}}}$.
@@
Moreover, the construction (3.6.3) can be twisted by $t$, yielding a definable map
$\widetilde{{\mathbf{M}}}^{\lozenge,t}_{{\mathcal{E}},a}\longrightarrow I_{\hat{\mu}}\widetilde{{\mathbf{M}}}^{\square,t}_{G,a,{\mathcal{E}}}.$
This map is a bijection in every finite field of large enough residue characteristic by an argument similar to [23, Lemma 6.8], hence in every pseudo-finite field. ∎
@@
Given $)t\in{\mathbb{X}}^{*}(\widehat{\mathbb{T}}/Z(X,\widehat{G}$)), we have canonical maps
@@
We also have the map $\lambda_{a}\colon{\mathbb{X}}_{*}({\mathbb{T}}/Z(X,G))\to\pi_{0}(\widetilde{{\mathcal{P}}}^{\square}_{G,a})$. We write abusively also $t$ for the image of $t$ in $\pi_{0}(\widetilde{{\mathcal{P}}}^{\square}_{G,a})$ through this composition of maps. In particular, given $)t\in{\mathbb{X}}^{*}(\widehat{\mathbb{T}}/Z(X,\widehat{G}$)), we write $\widetilde{{\mathcal{M}}}^{\square,t}_{G,a}$ for the twist of $\widetilde{{\mathcal{M}}}^{\square}_{G,a}$ by the image of $t$ in $\pi_{0}(\widetilde{{\mathcal{P}}}^{\square}_{G,a})$.
@@
For every pseudo-finite field $K$, $a\in\widetilde{{\mathbf{A}}}^{{\mathrm{ani}}}(K)$, $)t\in{\mathbb{X}}^{*}(\widehat{\mathbb{T}}/Z(X,\widehat{G}$)) and $s\in{\mathbb{X}}^{*}({\mathbb{T}}/Z(X,G))$, we have
@@
$\chi_{\mathrm{Psf},\mathrm{rel}}(I_{\hat{\mu}}\widetilde{{\mathbf{M}}}_{\widehat{G},a,{\mathcal{E}}_{H}}^{\square,s})=\chi_{\mathrm{Psf},\mathrm{rel}}(\widetilde{{\mathbf{M}}}_{{\mathcal{E}},a}^{\lozenge,s}).$
@@
$\sum_{s,t}\chi_{\mathrm{Psf},\mathrm{rel}}(\widetilde{{\mathbf{M}}}_{{\mathcal{E}},a}^{\square,s})=\left\lvert\pi_{0}(\widetilde{{\mathbf{P}}}^{\square}_{G,a})\right\rvert\left\lvert\pi_{0}(\widetilde{{\mathcal{P}}}^{\square}_{\widehat{G},a})\right\rvert\chi_{\mathrm{Psf},\mathrm{rel}}(\widetilde{{\mathbf{M}}}_{{\mathcal{E}},a}^{\lozenge},\mathrm{stab}).$
From Lemma 6.10 in [23], $\dim(\widetilde{\mathbf{M}}_{G})-w(\kappa_{a})=r_{G}^{H}(D)$, so we now conclude using Lemma 4.4.3 and Theorem 4.8.2. ∎
@@ 5.3. Reduction to motivic integrals
The restrictions of ${\mathcal{M}}_{G}$ and ${\mathcal{P}}_{G}$ to $U_{a}$ are denoted by ${\mathcal{M}}_{G,U_{a}}$ and ${\mathcal{P}}_{G,U_{a}}$. Essentially by definition of Henselization, see for example [37, Theorem I.4.2], we have
@@
For $t\in\pi_{0}(\widetilde{{\mathcal{P}}}^{\square}_{G,a})$ we then get a smooth DM-stack $\widetilde{{\mathcal{M}}}^{\square,t}_{G,U_{a}}$ and a corresponding definable $\widetilde{{\mathbf{M}}}^{\square,t}_{G,B_{a}}$.
@@
Let $K$ be a pseudo-finite field $K$. For every $a\in\widetilde{{\mathbf{A}}}^{{\mathrm{ani}}}(K)$, every $)t\in{\mathbb{X}}^{*}(\widehat{\mathbb{T}}/Z(X,\widehat{G}$)) and every $s\in{\mathbb{X}}^{*}({\mathbb{T}}/Z(X,G))$, we have